40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 15 17:19:09 2020, gautam, Update, IOO, MC F spectrum  MCF_MCL.pdf
    Reply  Tue Sep 15 19:17:30 2020, rana, Update, IOO, MC F spectrum  
       Reply  Wed Sep 16 09:43:41 2020, gautam, Update, IOO, MC F spectrum  MCF_MCL.pdf
Message ID: 15573     Entry time: Tue Sep 15 17:19:09 2020     Reply to this: 15574
Author: gautam 
Type: Update 
Category: IOO 
Subject: MC F spectrum  

There was an abrupt change in the MC_F spectrum between August 4 and August 5, judging by the summary pages - the 1 and 3 Hz resonances are no longer visible in the spectrum. Possibly, this indicates some electronics failure on the MC servo board / whitening board, the CDS settings don't seem to have changed. There is no record of any activity in the elog around those dates that would explain such a change. I'll poke around at 1X2 to see if anything looks different.


Update 1740: I found that the MCL / MCF cables were disconnected. So since August 5, these channels were NOT recording any physical quantity. Because their inputs weren't terminated, I guess this isn't a clean measurement of the whitening + AA noise, but particularly for MC_F, I guess we could use more whitening (see Attachment #1). Probably also means that the wandering ~10-30Hz line in the spectrogram is a electronics feature. The connections have now been restored and things look nominal again.

Attachment 1: MCF_MCL.pdf  62 kB  Uploaded Tue Sep 15 18:43:11 2020  | Hide | Hide all
MCF_MCL.pdf
ELOG V3.1.3-